THE MICROWAVE SPECTRUM OF THE FLUOROBENZENE-NEON DIMER

\underline{ROBB J. WILSON}, ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, Ann Arbor, Michigan 48109-1055, U.S.A..

The microwave spectra of three isotopomers of fluorobenzene-neon (normal, ^{22}Ne and d-5) have been observed using a Fourier transform spectrometer. Twenty-two transitions were fit to eight constants. The dimer has a stacked structure with the neon over the fluorobenzene ring in the symmetry plane. The center of mass distance (R_{cm}) is 3.44(1) Å, with the neon sitting between the center of mass and the center of the ring. The rotational constants have been determined as A = 1926.421(2) MHz, B = 1645.236(3) MHz, and C = 1279.734(1) MHz. The electric dipole moment was measured to be |μ_{a}| = 1.15(3) D, |μ_{b}| = 1.12(3) D, |μ_{t}| = 1.61(5) D.